A parallel-plate capacitor is made by using two sq
Answer #1

a) capacitance of a capacitor is given by

C=\frac{\varepsilon _{o}A}{d}

plugging in the values we get

C =4.8 x 10-13f

b) charge

Q =CV

plugging in we get

Q = 5.79 x 10-9C

c) electric field E is given by

E= V/d

E= 12/4.7 10-3

E =255.3 N/c

d) C=\frac{\varepsilon _{o}A}{d}

C =Co/2 = 2.41 x 10-13f

charge remains same

electric field remans same.

e) reactance

Xc =1/wC

Xc = 3.3 x108 ohms

Irms = (V/Xc)/sqrt2

Irms =1.07 x 10-7A

Irms for second case

Irms = 0.54 x10-7A